OpenStudy (anonymous):

volume of a solid whose base is 1/9 *x^2 + y^2=1 and having a square cross section perpendicular to the x axis

OpenStudy (anonymous):

volume. of a solid. not area of a cross section

OpenStudy (experimentx):

find the area ... y = sqrt(1 - 1/9 x^2) area = 4 * integrate from 0 to 3 sqrt(1 - 1/9) dx volume = area * height

OpenStudy (anonymous):

why is it from 0 to 3 also why is it 4 times the integral? shouldn't it be the function inside the integral squared?

OpenStudy (experimentx):